OpenStudy (anonymous):

ALGEBRA PROBLEM! HELPPPPPP<3 3 bananas are to be selected from a group of 9. In how many ways can this be done?

OpenStudy (experimentx):

9C3

OpenStudy (anonymous):

wait can you please help explain it to me?

OpenStudy (anonymous):

because 27 isn't one of the options!

OpenStudy (experimentx):

while SELECTING use this rule, if there are 'n' choices for 'r' places, then there are nCr choices, and nCr = \(\frac{n!}{(n-r)!r!}\) So, \( ^9C_3 = \frac{9!}{(9-3)!3!} \)

OpenStudy (experimentx):

looks like get got bug here.

OpenStudy (anonymous):

how many ways can you choose the first one? and then how many to choose from for the second? and now how many to choose from for the last one? multiply those together

OpenStudy (anonymous):

can someone draw it for me?

OpenStudy (anonymous):

please?

OpenStudy (anonymous):

i think it's 84?

OpenStudy (anonymous):

I have never seen a drawing of this before, and have no idea how to draw it.

OpenStudy (accessdenied):

Here's a smaller scale example of having 5 choices to choose 2

OpenStudy (anonymous):

Is it 84 or no?

OpenStudy (accessdenied):

84 sounds correct to me

OpenStudy (experimentx):

the answer is 9*8*7/(3*2) = 84

OpenStudy (experimentx):

your numbers are too big to draw ... try selecting 2 people out of 4

OpenStudy (accessdenied):

Here's a 'picture' of it, if you're interested. It's probably not too helpful of a chart to look at, but I guess it does confirm 84 as the answer. :D